Shift summary for day shift 7/17

Jim Hirschauer (jhirsch@cuhep.Colorado.EDU)
Sat, 17 Jul 1999 14:32:41 -0600 (MDT)

Shift Summary for day shift, 17/July/99

On Shift: Bob Hsiung, Jim Hirschauer

Hours of beam delivered at all intensities: 4.0
Hours of beam delivered at nominal intensity: 4.0
Hours of E832 data (special runs not included): 3.25
Hours of special runs (laser scans, muon runs...): 1.0 - laser scan
and ped run

Accounting of lost time: Linac problems, Main Injector problems, Run
initialization problems

========================================================================
Runs ended this shift:
Run 13926 - 5.9M events - BEAM_832
Run 13927 - 4.1M events - BEAM_832
Run 13928 - 27.8K events - Laser Scan
Run 13929 - 3.9K events - KTEV_PED

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Started out as normal E832 datataking, but then the Main Injector
crapped out, followed by LINAC problems. A Laser Scan and Pedestal run
were performed. We were attempting to init BEAM_832 running and
experiencing difficulties, as beam finally came back after a 3.5 hour
disappearance. Ed updated constants for the CsI constants between
runs.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13926 7:00 Take over from Owl shift

7:15 Blue Ice was replaced around Alcohol Bubbler

13926 124 7:54 MI power supply problem. MCR says it will take at
least 10 minutes to fix. EB is paged to update
CsI constants.

13927 8:01 Beam back. CsI constants are updated.
We init run 13927.

8:08 Run finally commences, but beam disserts.

8:22 Beam back

8:25 MI abort, no beam.

13927 43 9:05 MCR corrected bad power supply, beam is back.
Spills 1-43 are beamless.

9:50 Elliot admonishes us to run kfrend. We acquiesce
and ktevana errors go away.

9:55 Onsplit coerces us to load a new tape. We are
helpless but to comply.
Tape QKNL12 was loaded in ktev2, tps3d2.

13927 120 10:50 Beam leaves: "Linac Problems est 20 min."

11:15 Blue Ice changed again.

11:40 Linac problem est 1 hr. Run 13927 ended

13928 12:00 We run a Laser Scan

13929 12:45 We do a Pedestal Run

13:10 Ped run done. Updated ADGN13929

13:35 Init Run 13930 and await beam.

13:40 We hang up at init for quite sometime. Finally,
kumq button goes red w/ many errors.

drt_fresh_start, but hang up again, before kumq
have a chance to go bad, w/ red Etot button

drt_fresh_start again: kumq not red but Init times
out

14:15 Beam returns, Init problems continue...slowly.

14:40 Run is initialized after kumquats are manually
reset and reloaded w/ map-gen.

========================================================================
Follow up Notes on Previous Shift

o E/P (csigain) of CsI 3114 looks good now.
916 E/P is fine: RMS ~1.8%
o CsI constants has been updated before run 13927.

========================================================================
Items to Follow up on Next Shift

o keep collecting BEAM_832 data

o look into dead bit of AZLAT in ACC trigger only